UT Sep 05, 2019 - Restart Night #4


OSA: S. Allanson

ISp: E. Solheid

Mtn Manager: P. Hartley

ISA: B. Rothberg, O. Kuhn (Tucson)

Techincal Experts: J. Hill, J. Power (Tucson)

Software: P. Kubanek (Tucson)


-AGW-2 WFS is offline, P. Kubanek suspects the Leach controller is not working - will proceed without off-axis WFS on LUCI-2. See IT 7834.

Open for Skyflats

UT 02:03 - enclosure open - taking skyflats LUCI-Bino

UT 02:08 - LUCI-1 images are super dusty

-cube may not be parked on LUCI-1

-calling D. Miller to park cube (or verify LUCI-1 cube is parked)

-lots of dust at the top of LUCI-1 and LUCI-2

IT 7835 - dust doughnuts from LUCI window

luci1.20190905:

0001 - blind dark

0002-0030 skyflats K,Ks,H, J

luci2.20190905:

0002-0030 skyflats K,Ks,H, J

Breakdown:

0002-0012 - K

0013-0017 - Ks

0018-0023 - H

0024-0030 - J

UT 02:31 - restarting GCS - will start work on AGW-1 alignment

https://wiki.lbto.org/Commissioning/BGTechProcedures#Overview_of_AGW_45_Instrument_45_Telescope_Alignment

Taking test images at K and J of parked cubes - D. Miller confirms they are now parked (not sure if cubes were parked while taking flats earlier). (from tomorrow's email thread: AGW1 has cube on the stage with new mounting scheme. AGW2 has no cube mounted.)

0031 - K-band

0032 - J-band

-only improvement seems to be base of cube stage - less pronounced on LUCI-1, but stage for cube itself has ~same height on both sides y+170 pixels

Alignment of Guiding Hotspot for WFS

UT 02:42 Restart GCSL to input 2.54 arcsec hotspot params in LUCI_L.cfg (guiding hotspot was previously set to lab values of 249,211

UT 02:45 bino ACQUIRE preset on M5 star.

UT 02:51 left ACTIVE preset off-axis - star is in pinhole, now centering it with Zernikes blocked. -Y hotspot moves star up in lenslet. (249,200 is closer)

UT 03:02 on-axis ACTIVE preset - pupil center at 383,239

Petr patches GCS to have DX wfs in simulation while it won't communicate.

UT 03:09 Back to M5 star with both GCS restarted. GCSL failed on wfs due to patch.

UT 03:12 SX running regular GCS, DX running patch. Now stars on both sides, but one wasn't the M5 star.

UT 03:24 bino on-axis ACTIVE preset to Persson star (GUIDE on right side with Steve doing eyeball active optics)

UT 03:28 GCSL is complaining about not getting WFS images then stops. Hotspot 249,196 Pupil center (hotspot) 380,239 (IDL seems to want 386.7,238.7 but that doesn't look correct to my eye.) AGw1 startup complete with guiding hotspot andmutt pupil alignment complete.

UT 03:47 Restart GCSL (twice to avoid wfs problem?) and resend on-axis preset.

AGw probe alignment to rotator center (both sides)

(following the steps in BGTechProcedures)

stop guiding

move rotator

ptincrement

on OBS1 1000 setxy not found on Fedora

UT 03:56 90000 ms exposure /nfs/192.168.39.20/data/share/tcs/telescope.20190905/left_guider000001.fits , but nothing found at that path.

on TCS1 /lbt/data/config/tcs/GCS/left_guider000001.fits but no ds9 available on TCS1

UT 04:02 90000 ms /lbt/data/config/tcs/GCS/left_guider000002.fits
SX 110,101 220,202 235,241 need +14,-45 pixels decrease ROT homeoffset (14 pix = 0.7 arcsec = 0.4 mm = 26 counts) increase RAD ooffset (45 pix = 2.2 arcsec = 1320 microns)

DX 126,100 252,200 267,239 (hotspot 275,243)

rdwrconfig on both sides, then restart GCSs, resend presets. DX wants 0.17 0.82 mm ---> -11 -820

SX OBS2 /nfs/192.168.39.20/data/share/tcs/lbto.20190905/left_guider000001.fits No FILE AT THIS PATH ON OBS2 ?? /lbt/data/share/tcs/lbto.20190905

SX more like a saturated blob than a circle.
SX 118,81 236,162 251,201 (where hotspot is 249,196) = GOOD

DX 124,113 248,226 263,245 need +13,+2 pixels +0.65,+0.1 arcsec 0.28,0.06 mm -18 counts homeoffset (because John made a typo while entering the previous correction)

UT 04:49 Petr restarting AZCAMSERVER

Adjust SX M3 (part 1)

IDL can't deal with the wrong path in gcsclient return. Petr makes a patch of gcsclient during the following MOS error

Rotator ZP LUCIs

M52 astrometric mask

UT 05:26 - LUCI-2 MOS error - checking

El 53, Rotator 128

-LUCI grabber - motor steps in the grabber - didn't reach enough steps, didn't reach positive limit

-error while trying to grab the mask in storage (slot 22)

UT 05:50 -J. Power recovered LUCI-2 MOS error - manually moving astrometric mask to FPU

(footnote: J. Power made a joke about a MOS error so this is clearly her fault)

LUCI-1:

x-shift 4.84" y-shift -1.14" theta = 0.067

LUCI-2:

x-shift 3.61" y-shift -2.28" theta = -0.260

LUCI confirmation images look good

FM4 looks offset again on LUCI-2

Adjust SX M3 (part 2)

UT 06:12 preset back to Persson star (a different one)

DX AdSec rips

working in /lbto/data/20190905_lfbg/

stop wfsing on SX

pupil_collect, retval, exp=10.0
% Compiled module: STDEV.
% Compiled module: CIRCLE_FIT.
% Tag name MASK is undefined for structure WFSC_STRUCT.
% Execution halted at: CIRCLE_FIT         40 /home/lbto/idl/wfsc/circle/circle_fit.pro
%                      PUPIL_OUTER_FIT   307 /home/lbto/idl/wfsc/pupil/pupil_outer_fit.pro
%                      COLLIMATE_OUTPUT  132 /home/lbto/idl/wfsc/collimate/collimate_output.pro
%                      WFSC_PROCESS_FILE  174 /home/lbto/idl/wfsc/wfsc/wfsc_process_file.pro
%                      COLLIMATE_ONCE     62 /home/lbto/idl/wfsc/collimate/collimate_once.pro
%                      PUPIL_COLLECT      70 /home/lbto/idl/wfsc/pupil/pupil_collect.pro
%                      $MAIN$          

(Doug fixed this IDL problem the following evening.)


Imaging tests - Bino-LUCI

UT 06:28 - Moving to NGC 7027 for imaging and filter tests

-MOS error on LUCI-2

EL: 74

Rot: 238 - current (was slewing)

-pre-condition violation for FPU to storage, looks like at least one strain gauge on (maybe both on)

-zerored strain gauges, inti MOS, sending preset again - think non-linearity is in the FPU (mask 15, but it was mask 22 in there)

-stopping LUCI software

UT 06:39 - restarting LUCI software

called Igor as this is a weird software problem - the MOS switchbox shows mask 22 open

UT 12:15 - I. Sola suggests stopping and restarting all of the LUCI software - doing that now

-still comes up with mask #15 in FPU instead of #22

UT 07:27 - reconfiguring for LUCI-1 only, proceeding with night

IT 7837 - Critical IT for LUCI-2

UT 08:05 - proceeding with NGC 7027 as test of LUCI-1 (4pt dither cycle through all filters) - cycled through all except (

Phot/Spec-Phot Std LUCI-1

UT 08:45 - LUCI-1 phot standards

P290-D - cycle through z,Ks,K,H,J - Airmass 1.024 -seems to have a close companion, maybe not good to use again

L1:

0086-0089 z

0090-0093 Ks

0094-0097 K

0098-0101 H

0102-0105 J

UT 09:15 - moving to G191B2b - SpecPhot Std cycle thru z,Ks,K,H,J - Airmass 1.499

L1:

0106-0109 z

0110-0113 Ks

0114-0117 K

0118-0122 H

0123-0126 J

UT 09:50 - completed G191b2b

MIXED MODE LUCI/LBC Test

Reconfiguring for LUCI-1/LBC-Red to test mixed mode for upcoming AZ night

UT 10:11 - running dohybrid for LBC-Red (configured for LBC-R only first to collimate)

UT 10:21 - sending binocular LBC-R and LUCI-1 script (LBC script hand edited to send sync flag)

-moving to M76

UT 10:23 - resending scripts - LUCI-1 had grabbed the wrong guidestar

L1 0126 - Ks

-can see it in both

-seems to be successful

-LUCI dithering ok, first LBC dither was successful

UT 10:34 - LUCI-1 filter failure - Unknown FW2 - failure was when it was in Clear and for some reason attempted to move to Clear

-thinks its in the Blind, Switch 8 is on, 7,9,10 are off

-taking test exposure to see if actuall in blind - L1 0131 -is test

-Switch 7 on, 8,9,10 off - clear, - moved 194 steps in the right direction

-J.Power backtracing through log to verify the steps and the filters match

-IT on LUCI-1 FW2 -> the failure seemed to be when moving from Ks to Clear (switching in script from Ks to Br Gamma), it moved to Free06 then tried Free05 but failed.

Switches on F05 is: 0,1,1,0 - swtches for Blind is 0,1,0,0 (switch, 7, 8, 8, 10) - IF switch 9 was flaky, when the FW was initiailzed and switch 9 was misread as "0" value then the config would think its in the blind configuration

Concluded for collimation on both instruments - would want to send a sync preset - pick a field where there is both plenty of stars for LBC and an off-axis star for LUCI

UT 11:58 - finished mixed mode test


Back To SpecPhtStd

UT 12:00 - canceling preset (still in bino mode) - authorizing LUCI-1 only G191b2 - Airmass 1.104

L1: -skipping around script to change order because we are in twilight

0169-0172 - z

0173-0175 - Ks

0176-0178 - J

0179-0181 - H

0182-0184 - K

UT 12:30 - reauthorizing for Bino LUCI-1/LBC-R will attempt sky flats

-taking narrow band filter sky flats with LUCI-1

skyflat details:

Skyflats with luci1. Getting bright! Taken on blank field 03+31 (NDIT1)
  • luci1.20190905.0187.fits Blank03+31 2.51 clear Ks 16K
  • luci1.20190905.0188.fits Blank03+31 2.51 clear Ks 18K
  • luci1.20190905.0189.fits Blank03+31 2.51 clear Ks 20K
  • luci1.20190905.0190.fits Blank03+31 2.51 clear Ks 22K
  • luci1.20190905.0191.fits Blank03+31 2.51 Br_gam clear 3K
  • luci1.20190905.0192.fits Blank03+31 2.51 Br_gam clear 3K
  • luci1.20190905.0193.fits Blank03+31 6 Br_gam clear 10K
  • luci1.20190905.0194.fits Blank03+31 6 Br_gam clear 10K
  • luci1.20190905.0195.fits Blank03+31 6 Br_gam clear 10K
  • luci1.20190905.0196.fits Blank03+31 6 Br_gam clear 10K
  • luci1.20190905.0197.fits Blank03+31 2.51 FeII clear 15K
  • luci1.20190905.0198.fits Blank03+31 2.51 FeII clear 16K
  • luci1.20190905.0199.fits Blank03+31 2.51 FeII clear 16K
  • luci1.20190905.0200.fits Blank03+31 2.51 FeII clear 17K
  • luci1.20190905.0201.fits Blank03+31 2.51 FeII clear 18K

UT 12:50 - closing up

-Going to take imaging flats for LUCI-1

L1:

0202-0206 - z lights off

0207-0211 - z lights on

-stopping after z-band

UT 13:21 - putting LUCI-1 in safe mode

UPDATE: IT 7837 no longer critical - D. Thompson ran initialize (after marking slot 15 valid) and the correct mask appeared - he will troubleshoot and see if he can repeat this "bug"

HeI on LUCI-1 may be showing a light leak like seen earlier in J1 IT 6142

15 ITs tonight so far....


  • AGW1 and AGW2 checkout and alignment at LFBG (critical)
    • DONE initial wake-up
    • DONE wfs alignment (adjust guiding hotspot in twilight, measure pupil position) - AGW-2 offline for WFS
    • DONE align probe to put hotspot on rotator center
    • adjust M3 to minimize pupil wobble (critical)
    • refine wfs alignment
  • measure LFBG/RFBG field aberrations (high)
  • PEPSIPFU wake-up checks (high)
    • M3 alignment for PFU pupil (1 hr)
  • LUCI-1 and LUCI-2
    • DONE Rotation Zeropoint (Astrometric masks -required) - high (1hr)
    • Image Pointing and Acquisition Check - high (1 hr)
    • AGW-1 and AGW-2 Transform - high (2x2hrs - requires initial data, then followup to confirm transform)
    • Longslit check (dither along slit, offsets and dithering image tests) - high (1 hr)
    • non-sidereal check imaging - high (0.5hr)
    • non-sidereal check longslit - low (0.5 hr)
      • Target is asteroid (26) Prosperina which is up at the end of the night.
      • This is a preferred target since its spectra will serve as a standard that will benefit science programs.
      • May require 1.0 hr (not just 0.5 hr)
    • Twilight flats zJHKKs (narrow band if time) - low (2 hrs, twilght 1hr each)
      • DONE JHKKs; narrow band (LUCI-1 only)
    • Test Science OB -
      • DONE imaging - medium (1 hr)
      • longslit - medium (0.5 hr)
    • Photometric Standards/SpecStd
      • DONE LUCI-1
      • LUCI-2
  • DONE Mixed-mode LBC-R/LUCI-1 tests

-- BarryRothberg - 04 Sep 2019
Topic revision: r12 - 10 Sep 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback